# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v2

@@ 8  Computing the functions
For this, we use the sequence $\sigma$ of Definition˜1 to traverse $\partial P$, maintaining the visibility core in a Sliding-Window data structure (3). From here on, we fix the sequence $\sigma$ as the conforming sliding sequence computed by Theorem˜18. Notably, we only need to concern ourselves with $u$, where $\textnormal{{next}}(u)$ is realized only by ugly dominators (➷). Central to our analysis of the construction of this function are multiple interwoven charging arguments. In particular, we will charge almost all queries to the data structure to vertices of the shortest path from $u$ to $\textnormal{{next}}(u)$, and vertices of the visibility core itself.
@@ 8.1  Partitioning the function domain
Suppose $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷) $(g,[u,\textnormal{{next}}(u)])$. First, if $u=u_{h}$, then $u_{h}\in(i-1,i)$ and then by Lemma˜3, $\overline{h\,u_{h}}$ has a reflex-vertex in its interior. But then $h$ is the last intersection point of the supporting line of the first edge of the shortest path $S(u,j+1)$, and thus $g=h$, and in particular $v_{h}=\textnormal{{next}}(u)$. Thus $\textnormal{{next}}(u)$ is realized by a type ($\star$) dominator instead. Hence $(i)$ holds.
@@
Next, by definition, $g$ is the second intersection point of the supporting line $\ell_{u}$ of the first edge of the shortest path $S(u,j+1)$ with $\mathcal{E}[i-1,j+1]$. As $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷) $g$ is not a vertex of $\mathcal{E}[i-1,j+1]$. Let $C$ be the counter-clockwise chain of the boundary of $\mathcal{E}[i-1,j+1]$ that is right of $\ell_{u}$. Let $h^{\prime}$ be the last vertex in $C$. Let $u^{\prime},v^{\prime}]$ be the maximal visibility of $h^{\prime}$ in [i1,j+1][i-1,j+1]. This is unique, as $h^{\prime}\in\mathcal{E}[i-1,j+1]$ and $S(u,\textnormal{{next}}(u))$ is left turning (by Lemma˜3). As $h^{\prime}$ is right of $\ell_{u}$, it can see $u$, so $u^{\prime}\leq u$. Furthermore, by maximality of $\textnormal{{next}}(u)$, and the fact that $h^{\prime}$ induces a bad dominator ($\star$) whose visibility contains $u^{\prime},v^{\prime}]$, it must be that $g$ sees further than $v^{\prime}$. It follows that the point $g$ lies right of the first edge of the shortest path $S(h^{\prime},j+1)$, which coincides with the supporting line of $\overline{h^{\prime}\,v^{\prime}}$. Let $x^{\prime}$ be the vertex of $S(h^{\prime},j+1)$ defining this first edge. Any other vertex of $C$, and any other vertex of $\mathcal{E}[i-1,j+1]$ that sees $u$, is left of $\overline{h^{\prime}\,x^{\prime}}$ and thus has $x^{\prime}$ on its shortest path to $j+1$. Thus it cannot see $v^{\prime}$. Further, the guard $g$ must be on the edge after $h^{\prime}$ in counter-clockwise order. As $g$ is below the supporting line of $\overline{h^{\prime}\,v^{\prime}}$, so is the entirety of the edge $e$. Thus $h^{\prime}$ sees its two defining edges and hence is in $\mathcal{B}_{\sigma}(i,j)$, and in particular, $h^{\prime}=h$, and $u^{\prime},v^{\prime}]=[u_{h},v_{h}]$, and $g$ is on $e$, and $(iii)$ holds. By general position, $x^{\prime}$ is not colinear with $e$, and thus we have that both $u^{\prime}=u_{h}$ and $v^{\prime}=v_{h}$ lie strictly left of $e$, and thus $(ii)$ holds.
@@
Now, conversely, assume $(i)$, $(ii)$, and $(iii)$ hold. Then (i) implies that $u$ can see a point $g^{\prime}$ on the edge $e$. By $(iii)$, this guard $g^{\prime}$ sees further than $v_{h}$. But then it must be dominated by an ugly dominator (➷). In particular, the next vertex (in counter-clockwise order) after $h$ of $\mathcal{E}[i-1,j+1]$ can see the furthest among all guards on the edge $e$, and in particular further than $h$. Thus, $u$ does not see this next vertex, and the supporting line of the first edge of the shortest path $S(u,j+1)$ intersects the edge $e$, defining the the ugly dominator (➷) $g$ of $u$, concluding the proof. ∎
@@
Lemma 26.
The sets $\{I_{i,j,e}|(i,j,e)\in\mathcal{T}\}$ can be computed in $O(n\log n)$ time given the set $\mathcal{B}_{\sigma}$.
Proof.
Let us first consider the partition of $\partial P$ into sets $C_{i,j}=\{u\in\partial P|u\in[i-1,i),\textnormal{{next}}(u)\in(j,j+1]\}$. Each $C_{i,j}$ is contiguous, and there are at most $O(n)$ non-empty such $C_{i,j}$. These can also be computed by Lemma˜9 and Lemma˜10 in $O(n\log n)$ time. Next, consider the refinement of the set $C_{i,j}$ via the arrangement of $\{[u_{h},v_{h}]|(h,[u_{h},v_{h}])\in\mathcal{B}_{\sigma}(i,j)\}$ into cells $C_{i,j,h}$. By Lemma˜10, each guard in $\mathcal{B}_{\sigma}$ is in a contiguous subsequence of $\{\mathcal{B}_{\sigma}(i,j)|(i,j)\in\sigma\}$, there are again only $O(n)$ non-empty sets among all such $C_{i,j,h}$, and they can be computed in $O(n\log n)$ time via $\mathcal{B}_{\sigma}$ and $C_{i,j}$. Finally, associate every $C_{i,j,h}$ with the guard $(h^{*},[u_{h},v_{h}])\in\mathcal{B}_{\sigma}(i,j)$ (or more precisely, the edge $e$ of $\mathcal{E}[i-1,j+1]$ that is after $h^{*}$ in counter-clockwise order) that maximizes $v_{h}$ subject to $C_{i,j,h}\subset[u_{h},v_{h}]$. This can be done via Theorem˜21. Now consider only those $C_{i,j,h}$ such that its associated guard and edge fulfill the conditions of Lemma˜25. The set of cells among them associated to the same guard $(h^{*},[u_{h},v_{h}])$ is contiguous, as $i\in[u_{h},v_{h}]$, and hence any later $C_{i,j^{\prime},h^{\prime}}$ is also contained in $u_{h},v_{h}]$. The union of these sets associated to $h^{*}$ and with it, associated to $e$ define the set $I_{i,j,e}$. By construction, there are at most $O(n)$ many of these which are non-empty, and these can be computed in $O(n\log n)$ time from $C_{i,j,h}$ via a simple linear scan. ∎
Definition 11.
@@
Proof.
Let $I_{i,j,e}=[s,t]\subset[i-1,i]$. Let $g_{s}$ be the guard on edge $e$ realizing $\textnormal{{next}}(s)$, and let $g_{t}$ be the guard on edge $e$ realizing $\textnormal{{next}}(t)$. For every $u\in[s,t]$, the realizing guard $g_{u}$ is in $\overline{g_{s}\,g_{t}}$. The left pivot of $u$ is the first vertex of the shortest path $S(g_{u},u)$ and in particular, by the convexity of S(i1,j+1)S(i-1,j+1) on the shortest path $S(g_{t},s)$. Conversely, for every vertex $x$ in the shortest path $S(g_{t},s)$ there is a $u$ such that $\ell_{u}=x$. Similarly, the set of right pivots for $u\in[s,t]$ is precisely the set of vertices in the shortest path $S(g_{s},\textnormal{{next}}(t))$. These can be computed in time $O(\log n+|\{x|L_{i,j,e,x}\neq\emptyset\}|+|\{y|R_{i,j,e,y\neq\emptyset}\}|)$ using ˜1. The non-empty sets $L_{i,j,e,x}$ and $R_{i,j,e,y}$ are contiguous, and hence the non-empty sets $C_{i,j,e,x,y}$ can be computed by a simple linear scan in time $O(|\{x|L_{i,j,e,x}\neq\emptyset\}|+|\{y|R_{i,j,e,y\neq\emptyset}\}|)$ concluding the proof. ∎
Lemma 29.
@@
Proof.
For $u\in C_{i,j,e,x,y}$, $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷). That is, its realizing guard $g$ is the intersection of the supporting line of $\overline{u\,x}$ with (the supporting line of) $e$. Similarly, given $g$, $\textnormal{{next}}(u)$ is the intersection of the supporting line of $\overline{g\,y}$ with (the supporting line of) $[j,j+1]$. Both of these functions are Möbius transformations, i.e., functions of the form $\frac{A+Bu}{C+Du}$, where $A$, $B$, $C$, and $D$ depend on the coordinates of the start- and end points of $[i,i+1]$, $e$, $[j,j+1]$, $x$ and $y$. As Möbius transformations are closed under concatenation, their concatenation is of the form $\frac{A+Bu}{C+Du}$ as well, concluding the proof. ∎
Thus, if in total only $O(n)$ of the sets $L_{i,j,e,x}$ and $R_{i,j,e,y}$ are non-empty, then we can compute a representation of the next function of $O(n)$ pieces. To this end, consider the set $\hat{\mathcal{U}}$ of points $u\in\partial P$ such that $\textnormal{{next}}(u)$ is realized by only bad guards. The realizing guard $(g,[u,\textnormal{{next}}(u)])$ has that $\sphericalangle(u,g,\textnormal{{next}}(u))>\pi$, and that $S(u,\textnormal{{next}}(u))$ is left-turning. Consider the subset $\mathcal{U}\subset\hat{\mathcal{U}}$ for which the shortest path $S(u,\textnormal{{next}}(u))$ has at least two inner vertices. By Lemma˜25 and Lemma˜3, $u\in\partial P$ such that neither $u$ nor $\textnormal{{next}}(u)$ is a vertex of $P$ and $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷) in $\mathcal{U}$. For a vertex $x$ of $P$, let $L_{x}\subset\mathcal{U}$ be the set of points $u\in U$ such that $x$ is the first inner vertex of the shortest path $\pi_{u}=S(u,\textnormal{{next}}(u))$. Let $L_{x}$ be the number of connected components of $L_{x}$ as a subset of $\mathcal{U}$. In the next section, we show that
@@ 8.2  Bounding the number of pivot events
We begin by showing a stronger version of Theorem˜7. For this, we first define a relation on the vertices of the polygon, where $x\prec y$, if $x$ appears before $y$ in $\pi_{u}=S(u,\textnormal{{next}}(u))$ for some $u\in\mathcal{U}$, which will induce a strict order.
Lemma 30.
@@
The loop $L$ is simple and clockwise. First, observe that $\overline{\pi_{i}}$ and $\overline{\pi_{i+1}}$ intersect in $x_{i+1}$ (or the intersection point between $x_{1}$ and $x_{m}$), and as $\overline{\pi_{i}}$ and $\overline{\pi_{i+1}}$ are both left-turning shortest paths they cannot intersect in any other point. It remains to show that $\overline{\pi_{i}}$ and $\overline{\pi_{j}}$ do not intersect if $j\not\in\{i-1,i,i+1\}$. Assume otherwise. Then either $x_{j}$ or $x_{j+1}$ lies in $P_{i}$ contradicting the clockwise order of $x_{i}$, $x_{i+1}$, $x_{j}$ and $x_{j+1}$. Thus, $L$ is a simple loop. Furthermore, as each $\pi_{i}$ is left-turning and contained in $P$, the loop $L$ must be clockwise.
Figure 15: Construction of three guards guarding all of $\partial P$ from the proof of Lemma˜30. Illustrated is the fact that $u_{i_{1}},\textnormal{{next}}(u_{i_{1}})]$ contains $p_{3}^{*},p_{2}^{*}]$, which implies that $g_{i_{1}}$ sees $p_{3}^{*},p_{2}^{*}]$.
@@
Let us for now suppose, there are three points $p_{1}$, $p_{2}$, and $p_{3}$ on $L$, with tangents $h_{1}$, $h_{2}$ and $h_{3}$ defined by $\ell_{1}$, $\ell_{2}$, and $\ell_{3}$, such that $(i)$ each $h_{i}$ contains all three points $p_{1}$, $p_{2}$, and $p_{3}$, $(ii)$ $h_{1}\cap h_{2}\cap h_{3}$ is bounded, and $(iii)$ for the three $\pi_{i_{1}}$, $\pi_{i_{2}}$ and $\pi_{i_{3}}$ containing $p_{1}$, $p_{2}$ and $p_{3}$, no two are equal. Then there are three guards guarding all of $\partial P$ (refer to Figure˜15): As $h_{1}\cap h_{2}\cap h_{3}$ is bounded, we can pair $p_{1}$ with a point $p_{1}^{*}\in\partial P$ that is in $h_{2}\cap h_{3}$, if $p_{1}$ is not already on $\partial P$, by shooting a ray in $P$ starting in $p_{1}$ away from the intersection point $\ell_{2}\cap\ell_{3}$. As $h_{1}$ contains the point $\ell_{2}\cap\ell_{3}$, and the ray cannot intersect $\pi_{i_{1}}$ a second time before hitting the polygon boundary, we have that $p_{1}^{*}\in P_{i_{1}}$, unless $\pi_{1}$ was already in $P_{i_{1}}$, in which case we define $p_{1}^{*}=p_{1}$. Similarly, we find for $p_{2}$ and $p_{3}$ the points $p_{2}^{*}$ and $p_{3}^{*}$. The points $p_{1}^{*}$, $p_{2}^{*}$, and $p_{3}^{*}$ appear along $\partial P$ in clockwise order. Finally, note that $p_{2}^{*}$ and $p_{3}^{*}$ are in $h_{1}$, and thus do not lie in $u_{i_{1}},\textnormal{{next}}(u_{i_{1}})]$, which implies that $p_{3}^{*},p_{2}^{*}]\subset[u_{i_{1}},\textnormal{{next}}(u_{i_{1}})]$. [^2] Recall that intervals between points on $\partial P$ are defined in counter-clockwise order. In particular, there is a contiguous guard $(g_{1},[p_{3}^{*},p_{2}^{*}])$. Similarly, there are guards $(g_{2},[p_{1}^{*},p_{3}^{*}])$ and $(g_{3},[p_{2}^{*},p_{1}^{*}])$, which concludes the proof.
@@
The points $p_{1}$, $p_{2}$, and $p_{3}$ such that $(i)$, $(ii)$, and $(iii)$ hold, exist. Observe that indeed any number of such points already induce a set of three such points. The only condition that might fail, by removing such points is $(ii)$, but as the intersection of tangents together form a convex, bounded set, there must also be a subset of just three such tangents whose intersection is bounded.
We interpret $L$ as a simple polygon oriented clockwise. We begin, by picking $p_{1}$ and $h_{1}$ arbitrarily such that $p_{1}$ is not the intersection point of two paths $\pi_{i}$ and $\pi_{i+1}$. Let $p_{2}$ result from $p_{1}$ by shooting a ray from $p_{1}$ orthogonal to $\ell_{1}$. Pick $h_{2}$ arbitrarily, such that $p_{1}$ is not on $h_{2}$. Without loss of generality, $\ell_{1}$ and $\ell_{2}$ are not parallel, as otherwise $\textnormal{{next}}(u_{i_{1}}),u_{i_{1}}]$ and $\textnormal{{next}}(u_{i_{2}}),u_{i_{2}}]$ are disjoint, and there is even a set of guards guarding all of $\partial P$ of size $2$. Assume without loss of generality that $\ell_{1}$ and $\ell_{2}$ intersect right of the line segment $\overline{p_{1}\,p_{2}}$. As $L$ is a simple loop, there is a chain $s,t]\subset L$ such that $s$ is on $\ell_{1}$, $t$ is on $\ell_{2}$, and all of $[s,t]$ is in $h_{1}\cap h_{2}$, see Figure 16. If there is a point in $[s,t]$ with a tangent parallel to $\overline{p_{1}\,p_{2}}$, then we can choose $p_{3}$ as this point, as all of $[s,t]$ is left of $\overline{p_{1}\,p_{2}}$. Now assume that there is no point in $[s,t]$ with a tangent parallel to $\overline{p_{1}\,p_{2}}$. Consider the function $f$ which maps $x\in[s,t]$ to the distance from its orthogonal projection on $\ell_{1}$ to $p_{1}$. The function $f$ can only have one local minimum in $[s,t]$, as otherwise there is a point between two such local minima with tangent parallel to $\overline{p_{1}\,p_{2}}$. If $f$ has its local minimum at either $s$ or $t$, the chain $[s,t]$ is left-turning and there must be a point $p_{3}$ in $[s,t]$ whose tangent includes $p_{1}$ and $p_{2}$, as $[s,t]$ does not intersect $\overline{p_{1}\,p_{2}}$. Let instead $m\in(s,t)$ be the point where this minimum is attained.

### v4

@@ 8 Computing the functions
For this, we use the sequence $\sigma$ of Definition˜3.1 to traverse $\partial P$, maintaining the visibility core in a Sliding-Window data structure (3). From here on, we fix the sequence $\sigma$ as the conforming sliding sequence computed by Theorem˜6.11. Notably, we only need to concern ourselves with $u$, where $\textnormal{{next}}(u)$ is realized only by ugly dominators (➷). Central to our analysis of the construction of this function are multiple interwoven charging arguments. In particular, we will charge almost all queries to the data structure to vertices of the shortest path from $u$ to $\textnormal{{next}}(u)$, and vertices of the visibility core itself.
@@ 8.1 Partitioning the function domain
Suppose $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷) $(g,[u,\textnormal{{next}}(u)])$. First, if $u=u_{h}$, then $u_{h}\in(i-1,i)$ and then by Lemma˜4.4, $\overline{h\,u_{h}}$ has a reflex-vertex in its interior. But then $h$ is the last intersection point of the supporting line of the first edge of the shortest path $S(u,j+1)$, and thus $g=h$, and in particular $v_{h}=\textnormal{{next}}(u)$. Thus $\textnormal{{next}}(u)$ is realized by a type ($\star$) dominator instead. Hence $(i)$ holds.
@@
Next, by definition, $g$ is the second intersection point of the supporting line $\ell_{u}$ of the first edge of the shortest path $S(u,j+1)$ with $\mathcal{E}[i-1,j+1]$. As $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷) $g$ is not a vertex of $\mathcal{E}[i-1,j+1]$. Let $C$ be the counter-clockwise chain of the boundary of $\mathcal{E}[i-1,j+1]$ that is right of $\ell_{u}$. Let $h^{\prime}$ be the last vertex in $C$. Let $u^{\prime},v^{\prime}]$ be the maximal visibility of $h^{\prime}$ in [i1,j+1][i-1,j+1]. This is unique, as $h^{\prime}\in\mathcal{E}[i-1,j+1]$ and $S(u,\textnormal{{next}}(u))$ is left turning (by Lemma˜4.4). As $h^{\prime}$ is right of $\ell_{u}$, it can see $u$, so $u^{\prime}\leq u$. Furthermore, by maximality of $\textnormal{{next}}(u)$, and the fact that $h^{\prime}$ induces a bad dominator ($\star$) whose visibility contains $u^{\prime},v^{\prime}]$, it must be that $g$ sees further than $v^{\prime}$. It follows that the point $g$ lies right of the first edge of the shortest path $S(h^{\prime},j+1)$, which coincides with the supporting line of $\overline{h^{\prime}\,v^{\prime}}$. Let $x^{\prime}$ be the vertex of $S(h^{\prime},j+1)$ defining this first edge. Any other vertex of $C$, and any other vertex of $\mathcal{E}[i-1,j+1]$ that sees $u$, is left of $\overline{h^{\prime}\,x^{\prime}}$ and thus has $x^{\prime}$ on its shortest path to $j+1$. Thus it cannot see $v^{\prime}$. Further, the guard $g$ must be on the edge after $h^{\prime}$ in counter-clockwise order. As $g$ is below the supporting line of $\overline{h^{\prime}\,v^{\prime}}$, so is the entirety of the edge $e$. Thus $h^{\prime}$ sees its two defining edges and hence is in $\mathcal{B}_{\sigma}(i,j)$, and in particular, $h^{\prime}=h$, and $u^{\prime},v^{\prime}]=[u_{h},v_{h}]$, and $g$ is on $e$, and $(iii)$ holds. By general position, $x^{\prime}$ is not colinear with $e$, and thus we have that both $u^{\prime}=u_{h}$ and $v^{\prime}=v_{h}$ lie strictly left of $e$, and thus $(ii)$ holds.
@@
Now, conversely, assume $(i)$, $(ii)$, and $(iii)$ hold. Then (i) implies that $u$ can see a point $g^{\prime}$ on the edge $e$. By $(iii)$, this guard $g^{\prime}$ sees further than $v_{h}$. But then it must be dominated by an ugly dominator (➷). In particular, the next vertex (in counter-clockwise order) after $h$ of $\mathcal{E}[i-1,j+1]$ can see the furthest among all guards on the edge $e$, and in particular further than $h$. Thus, $u$ does not see this next vertex, and the supporting line of the first edge of the shortest path $S(u,j+1)$ intersects the edge $e$, defining the the ugly dominator (➷) $g$ of $u$, concluding the proof.
@@
Lemma 8.3.
The sets $\{I_{i,j,e}|(i,j,e)\in\mathcal{T}\}$ can be computed in $O(n\log n)$ time given $\mathcal{B}_{\sigma}$.
Proof 8.4.
Let us first consider the partition of $\partial P$ into sets $C_{i,j}=\{u\in\partial P|u\in[i-1,i),\textnormal{{next}}(u)\in(j,j+1]\}$. Each $C_{i,j}$ is contiguous, and there are at most $O(n)$ non-empty such $C_{i,j}$. These can also be computed by Lemma˜5.2 and Lemma˜5.3 in $O(n\log n)$ time. Next, consider the refinement of the set $C_{i,j}$ via the arrangement of $\{[u_{h},v_{h}]|(h,[u_{h},v_{h}])\in\mathcal{B}_{\sigma}(i,j)\}$ into cells $C_{i,j,h}$. By Lemma˜5.3, each guard in $\mathcal{B}_{\sigma}$ is in a contiguous subsequence of $\{\mathcal{B}_{\sigma}(i,j)|(i,j)\in\sigma\}$, there are again only $O(n)$ non-empty sets among all such $C_{i,j,h}$, and they can be computed in $O(n\log n)$ time via $\mathcal{B}_{\sigma}$ and $C_{i,j}$. Finally, associate every $C_{i,j,h}$ with the guard $(h^{*},[u_{h},v_{h}])\in\mathcal{B}_{\sigma}(i,j)$ (or more precisely, the edge $e$ of $\mathcal{E}[i-1,j+1]$ that is after $h^{*}$ in counter-clockwise order) that maximizes $v_{h}$ subject to $C_{i,j,h}\subseteq[u_{h},v_{h}]$. This can be done via Theorem˜6.19. Now consider only those $C_{i,j,h}$ such that its associated guard and edge fulfill the conditions of Lemma˜8.1. The set of cells among them associated to the same guard $(h^{*},[u_{h},v_{h}])$ is contiguous, as $i\in[u_{h},v_{h}]$, and hence any later $C_{i,j^{\prime},h^{\prime}}$ is also contained in $u_{h},v_{h}]$. The union of these sets associated to $h^{*}$ and with it, associated to $e$ define the set $I_{i,j,e}$. By construction, there are at most $O(n)$ many of these which are non-empty, and these can be computed in $O(n\log n)$ time from $C_{i,j,h}$ via a simple linear scan.
Definition 8.5.
@@
Proof 8.7.
Let $I_{i,j,e}=[s,t]\subseteq[i-1,i]$. Let $g_{s}$ be the guard on edge $e$ realizing $\textnormal{{next}}(s)$, and let $g_{t}$ be the guard on edge $e$ realizing $\textnormal{{next}}(t)$. For every $u\in[s,t]$, the realizing guard $g_{u}$ is in $\overline{g_{s}\,g_{t}}$. The left pivot of $u$ is the first vertex of the shortest path $S(g_{u},u)$ and in particular, by the convexity of S(i1,j+1)S(i-1,j+1) on the shortest path $S(g_{t},s)$. Conversely, for every vertex $x$ in the shortest path $S(g_{t},s)$ there is a $u$ such that $\ell_{u}=x$. Similarly, the set of right pivots for $u\in[s,t]$ is precisely the set of vertices in the shortest path $S(g_{s},\textnormal{{next}}(t))$. These can be computed in time $O(\log n+|\{x|L_{i,j,e,x}\neq\emptyset\}|+|\{y|R_{i,j,e,y\neq\emptyset}\}|)$ using ˜1. The non-empty sets $L_{i,j,e,x}$ and $R_{i,j,e,y}$ are contiguous, and hence the non-empty sets $C_{i,j,e,x,y}$ can be computed by a simple linear scan in time $O(|\{x|L_{i,j,e,x}\neq\emptyset\}|+|\{y|R_{i,j,e,y\neq\emptyset}\}|)$ concluding the proof.
Lemma 8.8.
@@
Proof 8.9.
For $u\in C_{i,j,e,x,y}$, $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷). That is, its realizing guard $g$ is the intersection of the supporting line of $\overline{u\,x}$ with (the supporting line of) $e$. Similarly, given $g$, $\textnormal{{next}}(u)$ is the intersection of the supporting line of $\overline{g\,y}$ with (the supporting line of) $[j,j+1]$. Both of these functions are Möbius transformations, i.e., functions of the form $\frac{A+Bu}{C+Du}$, where $A$, $B$, $C$, and $D$ depend on the coordinates of the start- and end points of $[i,i+1]$, $e$, $[j,j+1]$, $x$ and $y$. As Möbius transformations are closed under concatenation, their concatenation is of the form $\frac{A+Bu}{C+Du}$ as well, concluding the proof.
Thus, if in total only $O(n)$ of the sets $L_{i,j,e,x}$ and $R_{i,j,e,y}$ are non-empty, then we can compute a representation of the next function of $O(n)$ pieces. To this end, consider the set $\hat{\mathcal{U}}$ of points $u\in\partial P$ such that $\textnormal{{next}}(u)$ is realized by only bad guards. The realizing guard $(g,[u,\textnormal{{next}}(u)])$ has that $\sphericalangle(u,g,\textnormal{{next}}(u))>\pi$, and that $S(u,\textnormal{{next}}(u))$ is left-turning. Consider the subset $\mathcal{U}\subseteq\hat{\mathcal{U}}$ for which the shortest path $S(u,\textnormal{{next}}(u))$ has at least two inner vertices. By Lemma˜8.1 and Lemma˜4.4, $u\in\partial P$ such that neither $u$ nor $\textnormal{{next}}(u)$ is a vertex of $P$ and $\textnormal{{next}}(u)$ is realized by an ugly dominator (➷) in $\mathcal{U}$. For a vertex $x$ of $P$, let $L_{x}\subseteq\mathcal{U}$ be the set of points $u\in U$ such that $x$ is the first inner vertex of the shortest path $\pi_{u}=S(u,\textnormal{{next}}(u))$. Let $L_{x}$ be the number of connected components of $L_{x}$ as a subset of $\mathcal{U}$. In the next section, we show that
@@ 8.2 Bounding the number of pivot events
We begin by showing a stronger version of Theorem˜4.17. For this, we first define a relation on the vertices of the polygon, where $x\prec y$, if $x$ appears before $y$ in $\pi_{u}=S(u,\textnormal{{next}}(u))$ for some $u\in\mathcal{U}$, which will induce a strict order.
Lemma 8.10.
@@
[…]
Figure 15: Construction of three guards guarding all of $\partial P$ from the proof of Lemma˜8.10. Illustrated is the fact that $u_{i_{1}},\textnormal{{next}}(u_{i_{1}})]$ contains $p_{3}^{*},p_{2}^{*}]$, which implies that $g_{i_{1}}$ sees $p_{3}^{*},p_{2}^{*}]$.
@@
Let us for now suppose, there are three points $p_{1}$, $p_{2}$, and $p_{3}$ on $L$, with tangents $h_{1}$, $h_{2}$ and $h_{3}$ defined by $\ell_{1}$, $\ell_{2}$, and $\ell_{3}$, such that $(i)$ each $h_{i}$ contains all three points $p_{1}$, $p_{2}$, and $p_{3}$, $(ii)$ $h_{1}\cap h_{2}\cap h_{3}$ is bounded, and $(iii)$ for the three $\pi_{i_{1}}$, $\pi_{i_{2}}$ and $\pi_{i_{3}}$ containing $p_{1}$, $p_{2}$ and $p_{3}$, no two are equal. Then there are three guards guarding all of $\partial P$ (refer to Figure˜15): As $h_{1}\cap h_{2}\cap h_{3}$ is bounded, we can pair $p_{1}$ with a point $p_{1}^{*}\in\partial P$ that is in $h_{2}\cap h_{3}$, if $p_{1}$ is not already on $\partial P$, by shooting a ray in $P$ starting in $p_{1}$ away from the intersection point $\ell_{2}\cap\ell_{3}$. As $h_{1}$ contains the point $\ell_{2}\cap\ell_{3}$, and the ray cannot intersect $\pi_{i_{1}}$ a second time before hitting the polygon boundary, we have that $p_{1}^{*}\in P_{i_{1}}$, unless $\pi_{1}$ was already in $P_{i_{1}}$, in which case we define $p_{1}^{*}=p_{1}$. Similarly, we find for $p_{2}$ and $p_{3}$ the points $p_{2}^{*}$ and $p_{3}^{*}$. The points $p_{1}^{*}$, $p_{2}^{*}$, and $p_{3}^{*}$ appear along $\partial P$ in clockwise order. Finally, note that $p_{2}^{*}$ and $p_{3}^{*}$ are in $h_{1}$, and thus do not lie in $u_{i_{1}},\textnormal{{next}}(u_{i_{1}})]$, which implies that $p_{3}^{*},p_{2}^{*}]\subseteq[u_{i_{1}},\textnormal{{next}}(u_{i_{1}})]$. [^1] Recall that intervals between points on $\partial P$ are defined in counter-clockwise order. In particular, there is a contiguous guard $(g_{1},[p_{3}^{*},p_{2}^{*}])$. Similarly, there are guards $(g_{2},[p_{1}^{*},p_{3}^{*}])$ and $(g_{3},[p_{2}^{*},p_{1}^{*}])$, which concludes the proof.
@@
[…]
We interpret $L$ as a simple polygon oriented clockwise. We begin, by picking $p_{1}$ and $h_{1}$ arbitrarily such that $p_{1}$ is not the intersection point of two paths $\pi_{i}$ and $\pi_{i+1}$. Let $p_{2}$ result from $p_{1}$ by shooting a ray from $p_{1}$ orthogonal to $\ell_{1}$. Pick $h_{2}$ arbitrarily, such that $p_{1}$ is not on $h_{2}$. Without loss of generality, $\ell_{1}$ and $\ell_{2}$ are not parallel, as otherwise $\textnormal{{next}}(u_{i_{1}}),u_{i_{1}}]$ and $\textnormal{{next}}(u_{i_{2}}),u_{i_{2}}]$ are disjoint, and there is even a set of guards guarding all of $\partial P$ of size $2$. Assume without loss of generality that $\ell_{1}$ and $\ell_{2}$ intersect right of the line segment $\overline{p_{1}\,p_{2}}$. As $L$ is a simple loop, there is a chain $s,t]\subseteq L$ such that $s$ is on $\ell_{1}$, $t$ is on $\ell_{2}$, and all of $[s,t]$ is in $h_{1}\cap h_{2}$, see Figure 16. If there is a point in $[s,t]$ with a tangent parallel to $\overline{p_{1}\,p_{2}}$, then we can choose $p_{3}$ as this point, as all of $[s,t]$ is left of $\overline{p_{1}\,p_{2}}$. Now assume that there is no point in $[s,t]$ with a tangent parallel to $\overline{p_{1}\,p_{2}}$. Consider the function $f$ which maps $x\in[s,t]$ to the distance from its orthogonal projection on $\ell_{1}$ to $p_{1}$. The function $f$ can only have one local minimum in $[s,t]$, as otherwise there is a point between two such local minima with tangent parallel to $\overline{p_{1}\,p_{2}}$. If $f$ has its local minimum at either $s$ or $t$, the chain $[s,t]$ is left-turning and there must be a point $p_{3}$ in $[s,t]$ whose tangent includes $p_{1}$ and $p_{2}$, as $[s,t]$ does not intersect $\overline{p_{1}\,p_{2}}$. Let instead $m\in(s,t)$ be the point where this minimum is attained.